OpenStudy (anonymous):

Simplify 8(R + 6) - 2R: 6R + 6 6R + 46 or 6R + 48

rishavraj (rishavraj):

a(x + y) =ax + ay

OpenStudy (anonymous):

?

rishavraj (rishavraj):

whts 8(R + 6) ...???

OpenStudy (anonymous):

8(R-6)-2R=_____

rishavraj (rishavraj):

the questions is 8(R + 6) - 2R = ? so the hint is a(x + y) = ax +ay

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

srry for the confusion

rishavraj (rishavraj):

yw :))

OpenStudy (anonymous):

like my profile photo?i'm a big fan of power rangers

OpenStudy (anonymous):

the awnser is C

OpenStudy (anonymous):

the awnser is C